OpenStudy (zeus4263):

A box is to be constructed from a sheet of cardboard that is 20 cm by 50 cm by cutting out squares of length x by x from each corner and bending up the sides. What is the maximum volume this box could have? (Round your answer to two decimal places.)

OpenStudy (carolina4567):

So 0 < x < 10. That leaves us with x=4.40 as the solution. The size of the box: L=50-2x=50-8.80=41.20 W=20-2x=20-8.80=11.20 H=x=4.40 Maximum Volume of the box=(41.20)(11.20)(4.40)=2030.34 cm^3

OpenStudy (zeus4263):

Thank you!!!

OpenStudy (carolina4567):

No Problem :)